OpenStudy (anonymous):

in a right triangle find the length of the side not give b=2, c=sqrt 13

OpenStudy (anonymous):

a^2+b^2=c^2 a^2=c^2-b^2 =(sqrt13)^2-2^2 a^2=9 a=3 :)

OpenStudy (anonymous):

the key thing is you know that you have a right triangle so pythagorus can be apllied :)

OpenStudy (anonymous):

Thank you Mythias

OpenStudy (anonymous):

hey no prob any time :)

OpenStudy (anonymous):

@Mythias , make sure you don't give the answer directly. You direct them towards the answer. Read the Opensource guidelines